OpenStudy (anonymous):

a^5times a tiemsa^0 timesa^-3

OpenStudy (anonymous):

oops i missed an \(a\)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

\[a^5\times a\times a^0\times a^{-3}\] \[\large =a^{5+1+0+(-3)}\] \[=a^3\]